How does the nuclear mass of U-236 compare to the sum of the mass of the Ba-141, Kr-92 and 3 neutrons?
Answer #1

mass of U-236 = 236.046 amu  

( amu is atomic mass unit = 1.6737710-27 kg )

mass of Ba-141 = 140.914 amu

mass of Kr-92 = 91.926 amu

mass of neutron = 1.00866 amu

mass of ( Ba-141 + Kr-92 + 3 neutron ) = ( 140.914 +91.926+3.026 ) amu = 235.866 amu

mass difference between U-236 and ( Ba-141 + Kr-92 + 3 neutron ) is = ( 236.046 - 235.866 ) amu = 0.18 amu